If a retailer sells a coat for $139.99, what was his markup if he bought the coat for

    $80.00? (10 pts)

                                    Markup in actual dollars                                 ________

                                    Express markup as a % of cost                       ________

                                    Express markup as a % of selling price           ________
Answer #1

markup in dollars = selling price - cost

= 139.99 - 80

= 59.99

markup as cost = markup in dollars/cost

= 59.99/80

= 74.99%

markup as selling price = markup in dollars/selling price

= 59.99/139.99

= 42.85%